# Ultralight dark matter detection with trapped-ion interferometry

### v1

@@ Abstract
Introduction. Determining the microphysical nature of dark matter (DM) remains one of the outstanding problems in particle physics and cosmology. The parameter space of DM mass and coupling strengths to Standard Model (SM) particles is vast. Covering all possible candidates therefore requires a similarly broad experimental program.
@@
In this Letter, we propose searching for kinetically-mixed dark photon dark matter and ALPs with trapped-ion interferometers. Our concept takes advantage of the enormous improvement in the control and development of small-scale ion traps, primarily for quantum computing, and makes use of the associated advances in manipulation of the ion’s quantum states [58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. Leveraging the entanglement between the ion’s spin and motional degrees of freedom, trapped-ion matter-wave interferometers using “Schrödinger cat” (also known as “cat”) states [58], have been proposed to measure the Sagnac effect [63]. Here we demonstrate that a trapped-ion matter-wave interferometer using a cat state is sensitive to the Aharonov-Bohm phase induced by weakly-coupled dark matter fields. Considering the relevant boundary conditions, we demonstrate that such a quantum sensor can probe unexplored regions of ALP and DP parameter space.
Trapped-ion Interferometry. In this work, we consider the protocol proposed in Ref. [63]. An ion of mass $m_{\rm ion}$ is initially cooled and prepared in a linear superposition of internal quantum states $\ket{\uparrow}$ and $\ket{\downarrow}$ inside a harmonic and rotationally symmetric trap. The two wavepackets are subsequently set in counter-propagating orbital motion around the trap centre following $N$ successive spin-dependent kicks (SDKs), which entangle the motional and internal degrees of freedom and separate the two states by $2Nk_{\rm eff}$ in momentum space, and a non-adiabatic trap displacement $y_{d}$ in a direction orthogonal to the SDKs. As a result of their dynamics around the trap centre, the ion is in a linear superposition of spatially delocalised coherent states (i.e., cat states). After an integer number of revolutions, corresponding to an interrogation time $\Delta t$, an interference measurement is performed in the qubit basis {$\{\ket{+},\ket{-}\}$. A complete description of the protocol is given in the Supplemental Material. In the set-up envisaged in this work, we assume that the trap confines the motion of the ion to two dimensions, within which each wavepacket encloses an area $|\bm{\Sigma}|$ after a single revolution. Crucially, since each wavepacket has the same electric charge, the trapped ion will accumulate an Aharonov-Bohm phase shift that depends exclusively on the magnetic vector potential $\bm{A}$:
@@
The second line of Eq. (2) may equivalently be interpreted as the ion’s dynamical Zeeman phase shift in the presence of a background magnetic field, where the effective magnetic moment of the ion is defined as $\mu=2Nk_{\rm eff}\,y_{d}(m_{e}/m_{\mathrm{ion}})\mu_{B}$, where $m_{e}$ is the electron mass and $\mu_{B}$ is the Bohr magneton. For the parameters used in our projections, $\mu/\mu_{\mathrm{B}}\approx 1$ so that, after an interrogation time $\Delta t=1$ s, a detector employing a single ion limited by shot noise would have a sensitivity to a magnetic field of approximately $5\times 10^{-12}~{}\mathrm{T}/\sqrt{\mathrm{Hz}}$. We note that the noise level is significantly greater than that of other proposed dark matter experiments in this mass range [52, 55]. However, as shown below, the modest shielding we assume in our work means that the expected signal in our setup is much larger than was assumed in those works, and we find that the tradeoff between signal and noise that comes with having modest shielding of the apparatus is beneficial to the overall sensitivity (see Fig. 1).
Dark Matter Signals. Various dark matter models can generate small magnetic fields that are potentially detectable with the trapped-ion interferometer technique. The two of particular interest in the mass range to which the technique is most sensitive are kinetically-mixed dark photons and axion-like particles.
Dark Photon Dark Matter. Kinetically-mixed dark photon dark matter is defined by the Lagrangian $\mathcal{L}\supset\tfrac{\epsilon}{2}F_{\mu\nu}F^{\prime\mu\nu}+\tfrac{1}{2}m_{A^{\prime}}^{2}A^{\prime}_{\mu}A^{\prime\mu}$, where the first term contains the coupling to the SM photon, and the second endows the dark photon with a mass. The SM and dark photon system can be rotated to the so-called “interaction” basis, which has the benefit of making it clear that dark photons couple to the SM through EM fields generated by an effective current, $\bm{j}^{\prime}_{\rm eff}=-\epsilon m_{A^{\prime}}^{2}\bm{A^{\prime}}$. The dark photon vector potential has a Rayleigh-distributed amplitude, but is on average given approximately by
@@
For the mass range we consider here, the dark matter Compton wavelength $\sim 1/m_{A^{\prime}}$ is very long compared to all scales in the experiment, such as the size of the detector, the size of the shielded room hosting the detector, and even Earth’s radius $R_{\oplus}$. Therefore, we can work in the quasi-magnetostatic limit to estimate the size of the DM-generated magnetic field, as terms involving time derivatives such as $\partial_{t}\bm{E}$ will be sub-dominant relative to terms with spatial derivatives in Maxwell’s equations. Using Ampère’s law, we estimate the DP-induced magnetic field from ×$\nabla\times\bm{B}_{{}_{\rm DP}}\simeq\bm{j}_{\rm eff}$, where the operator $\nabla$ will set the spatial scale over which the field is present. We therefore expect a typical DM-induced magnetic field magnitude of $B_{{}_{\rm DP}}\sim-\epsilon m_{A^{\prime}}^{2}LA^{\prime}$. As we shall see, the scale $L$ that is appropriate for the trapped-ion interferometer is given by the Earth’s radius $R_{\oplus}$. If we had not accounted for boundary conditions and solved the full wave equation for $B_{{}_{\rm DP}}$, we might have estimated the DM-induced magnetic field as $B_{{}_{\rm DP}}\sim\epsilon\,m_{A^{\prime}}\bm{v}\times\bm{A^{\prime}}$. This naïve result is only valid in the regime when $m_{A^{\prime}}vR_{\oplus}\gg 1$, which is not the situation considered in this analysis. The full expression for the dark photon-induced magnetic field with the Earth as a boundary was computed in Ref. [68], and is reproduced in the Supplemental Material for completeness. Importantly, careful computation suggests that the dominant magnetic field component is tangential to Earth’s surface (i.e., the DP-sourced magnetic field has nonzero components only along the $\hat{\bm{\phi}}$ and $\hat{\bm{\theta}}$ directions).
Axion Dark Matter. For axion or ALP dark matter with a photon interaction, $\mathcal{L}\supset-\tfrac{1}{4}g_{a\gamma\gamma}aF_{\mu\nu}\tilde{F}^{\mu\nu}$, EM fields are generated by an effective current $\bm{j}_{\rm eff}^{a}\simeq-g_{a\gamma\gamma}\partial_{t}a\bm{B}_{0}$, to leading order in the DM velocity $v$. The axion field is given approximately by
where $\varphi$ is a random phase. From the form of the effective current, it is clear that a background magnetic field is required to induce a signal. The ion is trapped in an RF Paul trap characterised by a DC and AC potential, such that there is no significant background magnetic field. Therefore, to leading order in the DM velocity, there is no signal arising from the axion interacting with the trap EM fields. Instead, an effect may arise from the interaction of the axion with Earth’s magnetic field, which induces a signal magnetic field whose magnitude is approximately $B_{a}\sim(g_{a\gamma\gamma}a)(m_{a}R_{\oplus})B_{\oplus}$ in the $m_{a}R_{\oplus}\ll 1$ limit, where $B_{\oplus}\sim 0.5\,\text{G}$ is the typical magnitude of Earth’s magnetic field. The full axion-induced magnetic field resulting from coupling to Earth’s magnetic field was computed in Ref. [69], and is reproduced in the Supplemental Material for completeness. Crucially, as in the DP case, the magnetic field sourced by ALPs is tangential to the Earth’s surface.
Boundary Conditions Dictating Dark Matter Signals. The statements above about the magnitude of the expected DM-induced magnetic fields are contingent on applying the appropriate boundary conditions to the problem. A possible concern is that having a conductive boundary whose length scale $R$ is smaller than $R_{\oplus}$ might further reduce the signal fields by a factor $R/R_{\oplus}\ll 1$. However, the signals we are interested in measuring are typically in the frequency range $f\lesssim 10^{2}\,\text{Hz}$. Magnetic fields at such low frequencies are notoriously difficult to shield [70]. In a Paul trap, the ion is not fully enclosed by conductive metal. Therefore, there is magnetic field leakage into the region where the ion propagates. The laboratory hosting the ion trap might have a conductive boundary to reduce high-frequency EM noise. Computing the shielding efficiency of the precise geometry of the lab is beyond the scope of this Letter. However, intuition for the shielding efficiency can be developed by considering a shield with a greater degree of symmetry, such as a sphere. As described in detail in the Supplemental Material, the modulus of the shielding efficiency, defined as $\eta\equiv|\bm{B}_{\rm shield}/\bm{B}_{\rm naive}|$, depends on the sphere’s radius $R$, the thickness of the shield $\Delta$, the relative permeability of the shield $\mu$, and $\gamma\simeq(1+i)/\delta_{s}$, where $\delta_{s}=\sqrt{\omega\sigma\mu/2}$ is the skin depth. In the thin-shield limit, i.e., $\Delta\ll\delta_{s},R$, the modulus of the shielding efficiency is approximately
[…]
Noise Sources. A corollary of the poor shielding of low-frequency signals is that we anticipate ambient magnetic field noise to be the dominant limitation on the ion interferometer sensitivity for DM masses $m_{{}_{\rm DM}}\lesssim 10^{-14}\,\text{eV}$. The expected noise level at $f=1$ Hz is $B_{n}\simeq 10^{-11\pm 1}\,\text{T/}\sqrt{\text{Hz}}$ and scales as $1/f$ across the entire frequency range of interest [71, 72] [^1] Although measured noise levels in certain labs can be up to a factor of 10 higher at similar frequencies, see, e.g., [73]. These noise scalings appear in our sensitivity projections, with an order of magnitude error applied to account for temporal variations. Part of this ambient magnetic field noise could be due to vibrations of the apparatus, e.g., time-dependent changes in the orientation of $\hat{\bm{\Sigma}}$ relative to slow-varying ambient magnetic fields, which mimic a signal at higher frequencies. To mitigate this, one could conceivably run two versions of the experiment alongside one another, with only one being seismically isolated. A noise subtraction could then be performed by comparing the two datasets.
In order to achieve shot noise-limited sensitivity, various systematic effects associated to the trap itself must be controlled or mitigated. These have been discussed in detail in Ref. [64], so we only briefly summarise them here. Systematic effects typically affect either the phase, the enclosed area or the visibility. Such effects include differences in the spin-dependent kicks applied to the two ion states, spurious motion of the ion states in the trap, anharmonicities of the trap, and cross-coupling of the trap principal axes. Of these, it is control of the trap potential that is most crucial for ensuring the performance of the ion interferometer. For example, trap anharmonicities of the form $V(x,y)\supset\kappa_{x}x^{2}+\kappa_{y}y^{2}+\lambda_{x}x^{4}+\lambda_{y}y^{4}$ should be minimised. If $\lambda_{i}/\kappa_{i}\sim 0.1$, this would lead to a $10\,\%$ loss in visibility and a modification of the subtended area by the same amount. As the expected ratios $\lambda_{i}/\kappa_{i}\sim 10^{-4}$ [64], only minimal losses in visibility and changes in the interferometer area are expected. We provide more detailed discussion of the systematic effects in the Supplemental Material. In our sensitivity estimates, we assume they can be controlled at the level required to achieve shot noise-limited sensitivity.
Projected sensitivity. In Figure 1 we plot the projected power-averaged 95% upper limits on $\epsilon$ (left panel) and $g_{a\gamma\gamma}$ (right panel) for $T_{\mathrm{int}}=10^{8}$ s. As explained in detail in the Supplemental Material, the 95% upper limits on the couplings correspond to $\mathrm{SNR}\approx 1.6$ for $T_{\mathrm{int}}\gg\tau_{c}$ and $\mathrm{SNR}\approx 12.5$ for $T_{\mathrm{int}}\ll\tau_{c}$, where $\tau_{c}=2\pi/mv^{2}$ is the coherence time of the DM field. In purple, we show the projected benchmark sensitivity of a single-ion trap employing the parameters proposed in Ref. [63]. In particular, we envisage employing 171Yb+ and the (magnetically insensitive) internal ground hyperfine states |=|$\ket{\downarrow}=\ket{F=0,m_{F}=0}$ and |=|$\ket{\uparrow}=\ket{F=1,m_{F}=0}$ with frequency splitting $\omega_{a}/2\pi\approx 12.6$ GHz. The coherence time of the $\ket{\uparrow}$ state exceeds $10^{3}$ s [76], which allows for an interrogation time $\Delta t=1$ s. To drive transitions between qubit states, we assume the demonstrated two-photon stimulated Raman transitions using a laser at $355$ nm, which implies a momentum transfer $k_{\rm eff}=4\pi/355~{}\mathrm{nm}^{-1}$ to a wavepacket’s centre-of-mass per SDK [77]. In light of their high fidelity, we consider $N=100$ SDKs. Furthermore, we take the non-adiabatic displacement of the trap centre to be $y_{d}=100~{}\mu$m, which can be achieved by nano-second switching of the trapping potentials [78], and has recently been shown to yield displacements of $10~{}\mu$[79].
@@
Discussion. In this Letter we have studied the prospects of using trapped ions as matter-wave interferometers to search for electromagnetically-coupled ultralight dark matter, which imprints a non-zero Aharonov-Bohm-like phase on spin states of the ion.

### v2

@@ Abstract
I Introduction.
Determining the microphysical nature of dark matter (DM) remains one of the outstanding problems in particle physics and cosmology. The parameter space of DM mass and coupling strengths to Standard Model (SM) particles is vast. Covering all possible candidates therefore requires a similarly broad experimental program.
@@
In this work, we propose searching for kinetically-mixed dark photon dark matter and ALPs with trapped-ion interferometers. Our concept takes advantage of the enormous improvement in the control and development of small-scale ion traps, primarily for quantum computing, and makes use of the associated advances in manipulation of the ion’s quantum states [58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. Leveraging the entanglement between the ion’s spin and motional degrees of freedom, trapped-ion matter-wave interferometers using “Schrödinger cat” (also known as “cat”) states [58], have been proposed to measure the Sagnac effect [63], namely the relative phase shift between counter-propagating beams in a loop interferometer induced by the rotation of the interferometer’s reference frame [68]. Here we demonstrate that a trapped-ion matter-wave interferometer using a cat state is sensitive to the Aharonov-Bohm phase induced by weakly-coupled dark matter fields. Considering the relevant boundary conditions, we demonstrate that such a quantum sensor can probe unexplored regions of ALP and DP parameter space.
II Trapped-ion Interferometry.
[…]
@@
The second line of Eq. (2) may equivalently be interpreted as the ion’s dynamical Zeeman phase shift in the presence of a background magnetic field, where the effective magnetic moment of the ion is defined as $\mu=2Nk_{\rm eff}\,y_{d}(m_{e}/m_{\mathrm{ion}})\mu_{B}$, where $m_{e}$ is the electron mass and $\mu_{B}$ is the Bohr magneton. For the parameters used in our projections, $\mu/\mu_{\mathrm{B}}\approx 1$ so that, after an interrogation time $\Delta t=1$ s, a detector employing a single ion limited by shot noise (i.e., with phase shift sensitivity $1~\mathrm{rad}/\sqrt{\mathrm{Hz}}$) would have a sensitivity to a magnetic field of approximately $5\times 10^{-12}~\mathrm{T}/\sqrt{\mathrm{Hz}}$. We note that the noise level is significantly greater than that of other proposed dark matter experiments in this mass range [52, 55]. However, as shown below, the modest shielding we assume in our work means that the expected signal in our setup is much larger than was assumed in those works, and we find that the tradeoff between signal and noise that comes with having modest shielding of the apparatus is beneficial to the overall sensitivity (see Fig. 1).
III Dark Matter Signals
. Various dark matter models can generate small magnetic fields that are potentially detectable with the trapped-ion interferometer technique. The two of particular interest in the mass range to which the technique is most sensitive are kinetically-mixed dark photons and axion-like particles.
III.1 Dark Photon Dark Matter.
Kinetically-mixed dark photon dark matter is defined by the Lagrangian $\mathcal{L}\supset\tfrac{\epsilon}{2}F_{\mu\nu}F^{\prime\mu\nu}+\tfrac{1}{2}m_{A^{\prime}}^{2}A^{\prime}_{\mu}A^{\prime\mu}$, where the first term contains the coupling to the SM photon, and the second endows the dark photon with a mass. The SM and dark photon system can be rotated to the so-called “interaction” basis, which has the benefit of making it clear that dark photons couple to the SM through EM fields generated by an effective current, $\bm{j}^{\prime}_{\rm eff}=-\epsilon m_{A^{\prime}}^{2}\bm{A^{\prime}}$. The dark photon vector potential has a Rayleigh-distributed amplitude, but is on average given approximately by
@@ III Dark Matter Signals
where $\hat{\bm{n}}_{i}$, i=1, 2, 3i=1,\,2,\,3, is a set of orthonormal basis vectors, and $\varphi_{i}$ are random phases drawn from a flat distribution in $0,2\pi)$. The scalar potential $A_{0}^{\prime}\sim v|\bm{A^{\prime}}|$ is suppressed relative to the vector potential, since the Proca condition fixes $\partial_{\mu}A^{\prime\mu}=0$, and $|\bm{k}|\sim m_{A^{\prime}}v$ with $v\sim 10^{-3}\,c$ for virialised dark matter in the Milky Way.
For the mass range we consider here, the dark matter Compton wavelength $\sim 1/m_{A^{\prime}}$ is very long compared to all scales in the experiment, such as the size of the detector, the size of the shielded room hosting the detector, and even Earth’s radius $R_{\oplus}$. Therefore, we can work in the quasi-magnetostatic limit to estimate the size of the DM-generated magnetic field, as terms involving time derivatives such as $\partial_{t}\bm{E}$ will be sub-dominant relative to terms with spatial derivatives in Maxwell’s equations. Using Ampère’s law, we estimate the DP-induced magnetic field from $\nabla\times\bm{B}_{{}_{\rm DP}}\simeq\bm{j}_{\rm eff}$, where the operator $\nabla$ will set the spatial scale over which the field is present. We therefore expect a typical DM-induced magnetic field magnitude of $B_{{}_{\rm DP}}\sim-\epsilon m_{A^{\prime}}^{2}LA^{\prime}$. As we shall see, the scale $L$ that is appropriate for the trapped-ion interferometer is given by the Earth’s radius $R_{\oplus}$. If we had not accounted for boundary conditions and solved the full wave equation for $B_{{}_{\rm DP}}$, we might have estimated the DM-induced magnetic field as $B_{{}_{\rm DP}}\sim\epsilon\,m_{A^{\prime}}\bm{v}\times\bm{A^{\prime}}$. This naïve result is only valid in the regime when $m_{A^{\prime}}vR_{\oplus}\gg 1$, which is not the situation considered in this analysis. The full expression for the dark photon-induced magnetic field with the Earth as a boundary was computed in Ref. [69], and is reproduced in Appendix B for completeness. Importantly, careful computation suggests that the dominant magnetic field component is tangential to Earth’s surface (i.e., the DP-sourced magnetic field has nonzero components only along the $\hat{\bm{\phi}}$ and $\hat{\bm{\theta}}$ directions).
III.2 Axion Dark Matter.
For axion or ALP dark matter with a photon interaction, $\mathcal{L}\supset-\tfrac{1}{4}g_{a\gamma\gamma}aF_{\mu\nu}\tilde{F}^{\mu\nu}$, EM fields are generated by an effective current $\bm{j}_{\rm eff}^{a}\simeq-g_{a\gamma\gamma}\partial_{t}a\bm{B}_{0}$, to leading order in the DM velocity $v$. The axion field is given approximately by
where $\varphi$ is a random phase. From the form of the effective current, it is clear that a background magnetic field is required to induce a signal. The ion is trapped in an RF Paul trap characterised by a DC and AC potential, such that there is no significant background magnetic field. Therefore, to leading order in the DM velocity, there is no signal arising from the axion interacting with the trap EM fields. Instead, an effect may arise from the interaction of the axion with Earth’s magnetic field, which induces a signal magnetic field whose magnitude is approximately $B_{a}\sim(g_{a\gamma\gamma}a)(m_{a}R_{\oplus})B_{\oplus}$ in the $m_{a}R_{\oplus}\ll 1$ limit, where $B_{\oplus}\sim 0.5\,\text{G}$ is the typical magnitude of Earth’s magnetic field. The full axion-induced magnetic field resulting from coupling to Earth’s magnetic field was computed in Ref. [70], and is reproduced in Appendix B for completeness. Crucially, as in the DP case, the magnetic field sourced by ALPs is tangential to the Earth’s surface.
III.3 Boundary Conditions Dictating Dark Matter Signals.
The statements above about the magnitude of the expected DM-induced magnetic fields are contingent on applying the appropriate boundary conditions to the problem. A possible concern is that having a conductive boundary whose length scale $R$ is smaller than $R_{\oplus}$ might further reduce the signal fields by a factor $R/R_{\oplus}\ll 1$. However, the signals we are interested in measuring are typically in the frequency range $f\lesssim 10^{2}\,\text{Hz}$. Magnetic fields at such low frequencies are notoriously difficult to shield [71]. In a Paul trap, the ion is not fully enclosed by conductive metal. Therefore, there is magnetic field leakage into the region where the ion propagates. The laboratory hosting the ion trap might have a conductive boundary to reduce high-frequency EM noise. Computing the shielding efficiency of the precise geometry of the lab is beyond the scope of this work. However, intuition for the shielding efficiency can be developed by considering a shield with a greater degree of symmetry, such as a sphere. As described in detail in Appendix B.3, the modulus of the shielding efficiency, defined as $\eta\equiv|\bm{B}_{\rm shield}/\bm{B}_{\rm naive}|$, depends on the sphere’s radius $R$, the thickness of the shield $\Delta$, the relative permeability of the shield $\mu$, and $\gamma\simeq(1+i)/\delta_{s}$, where $\delta_{s}=\sqrt{2/\omega\sigma\mu}$ is the skin depth. In the thin-shield limit, i.e., $\Delta\ll\delta_{s},R$, the modulus of the shielding efficiency is approximately
For typical materials such as copper ($\mu=1,~\sigma=2.5\times 10^{7}\,\text{S/m}$) or ferromagnetic stainless steel ($\mu\sim 1000,~\sigma\sim 2\times 10^{4}\,\text{S/m}$), the shielding efficiency for a 1 mm thick layer of material in a 5 m radius “room” can be approximated using Eq. (5) for $\omega\lesssim 10^{-14}\,\text{eV}$ and is $\eta\gtrsim 0.9$. Therefore, unless the thickness of the shielding material is significant, for most of the parameter space we are studying, the shielding does not significantly reduce the signal magnetic field.
IV Noise Sources
. A corollary of the poor shielding of low-frequency signals is that we anticipate ambient magnetic field noise to be the dominant limitation on the ion interferometer sensitivity for DM masses $m_{{}_{\rm DM}}\lesssim 10^{-14}\,\text{eV}$. The expected noise level at $f=1$ Hz is $B_{n}\simeq 10^{-11\pm 1}\,\text{T/}\sqrt{\text{Hz}}$ and scales as $1/f$ across the entire frequency range of interest [72, 73] [^1] Although measured noise levels in certain labs can be up to a factor of 10 higher at similar frequencies, see, e.g., [74]. These noise scalings appear in our sensitivity projections, with an order of magnitude error applied to account for temporal variations. Part of this ambient magnetic field noise could be due to vibrations of the apparatus, e.g., time-dependent changes in the orientation of $\hat{\bm{\Sigma}}$ relative to slow-varying ambient magnetic fields, which mimic a signal at higher frequencies. To mitigate this, one could conceivably run two versions of the experiment alongside one another, with only one being seismically isolated. A noise subtraction could then be performed by comparing the two datasets.
In order to achieve shot noise-limited sensitivity, various systematic effects associated to the trap itself must be controlled or mitigated. These have been discussed in detail in Ref. [64], so we only briefly summarise them here. Systematic effects typically affect either the phase, the enclosed area or the visibility. Such effects include differences in the spin-dependent kicks applied to the two ion states, spurious motion of the ion states in the trap, anharmonicities of the trap, and cross-coupling of the trap principal axes. Of these, it is control of the trap potential that is most crucial for ensuring the performance of the ion interferometer. For example, trap anharmonicities of the form $V(x,y)\supset\kappa_{x}x^{2}+\kappa_{y}y^{2}+\lambda_{x}x^{4}+\lambda_{y}y^{4}$ should be minimised. If $\lambda_{i}/\kappa_{i}\sim 0.1$, this would lead to a $10\,\%$ loss in visibility and a modification of the subtended area by the same amount. As the expected ratios $\lambda_{i}/\kappa_{i}\sim 10^{-4}$ [64], only minimal losses in visibility and changes in the interferometer area are expected. We provide a more detailed discussion of systematic effects in Appendix A.4. In our sensitivity estimates, we assume they can be controlled at the level required to achieve shot noise-limited sensitivity.
V Projected sensitivity
@@
VI Discussion and conclusions
. In this work we have studied the prospects of using trapped ions as matter-wave interferometers to search for electromagnetically-coupled ultralight dark matter, which imprints a non-zero Aharonov-Bohm-like phase on spin states of the ion.